OpenStudy (anonymous):

MEDAL!! Alex is having wood flooring put in his rectangular living room.The area of the room is 234 feet.the length of the room is 18 feet.what is the width,in feet of the room?

OpenStudy (anonymous):

ok Area=l*w

OpenStudy (anonymous):

so 234=18*w

OpenStudy (anonymous):

so the width is 13 feet

OpenStudy (anonymous):

OHH ! thanks!! so I just needed to set up an equation

OpenStudy (anonymous):

ye ofcourse which says A=l*w

OpenStudy (anonymous):

cab you take a look at this last problems please??

OpenStudy (anonymous):

it's the very last problem

OpenStudy (anonymous):

which one the engineer?

OpenStudy (anonymous):

yeah that one

OpenStudy (anonymous):

why theres numbers over letter its messed up send it again

OpenStudy (anonymous):

is it showing good now??

OpenStudy (anonymous):

no but moreover the question makes no sense

OpenStudy (mikeymaximum):

@demonchild99

OpenStudy (mikeymaximum):

@LiteNing1337

OpenStudy (mikeymaximum):

@TheExpert

OpenStudy (mikeymaximum):

@TheRaggedyDoctor